# MauveSim: the instrument simulator software for the Mauve missionMauveSim: the instrument simulator software for the Mauve mission

### v2

@@ 4.3 Converting flux into counts
$\text{efficiency}=GR\cdot FB\cdot(1-CL)^{4}\cdot MT^{2}\cdot MS^{2}\cdot DCL\cdot WD\cdot OE\,.$ (12)
@@
[…]
@@ 4.4 Adding astrophysical and instrumental noise sources
Figure 7: Simulation of HD 143459 using an exposure time of 5 s. The final simulated observation and its components are shown as well. Panel (a): output Mauve spectrum as simulated by the software. Panel (b): net simulated signal source. Panel (c): photon noise. Panel (d): bias noise.
@@
Photon noise, which follows Poisson statistics, is modelled as the square root of the net signal in units of $e^{-}$. To simulate the distribution of photons striking the detector, we include a random component to the photon noise which is drawn from a Gaussian distribution with $\mu$=0 and $\sigma$=net $\sqrt{\text{net signal}[e^{-}]}$. Finally, to align the photon noise with the units of the source signal, photon noise is converted from $e^{-}$ to counts.
@@
where $T$ is the temperature of the detector. Based on this equation, the dark current is 445 counts/s at 21°C. This noise component is added by drawing random samples from a Poisson distribution with expectation value equal to $I_{\text{d}}(T)$ multiplied by the exposure time.
In addition, the bias noise is included. The bias is an offset applied to the ADC (Analogue-to-Digital Converter) to prevent negative counts during readout. The bias was measured in the lab and found to have an average value of 600 counts per pixel across the detector (Fig. 7 (d)).
The output of the MauveSim’s workflow is a 1D raw data spectrum similar to Fig. 7 (a). The simulator can also create a time series observation, where multiple frames obtained with the same exposure time are stacked up together to simulate a longer observation.
@@ 4.5 Signal-to-noise calculation
Figure 8: The S/N per pixel of the simulated data in Fig. 7.
Figure 9: S/N per resolution element on the simulation of HD 143459 for a exposure time of 5 s.
@@
where $S$ is the net signal in units of counts (Fig. 7 (b)); $\sigma_{S}$ is the photon noise ($\sqrt{S}$); $\sigma_{R}$ is the readout noise (equal to approximately 12 counts); $\sigma_{D}$ is the dark current noise (equal to the sqrt of the dark current, i.e. $\sqrt{\sim 2220\ \text{counts}}$ for the example in Fig. 7). Fig. 8 displays the S/N per pixel of HD 143459 for a 5 s integration.
@@ 5 Science case studies
Mauve’s performance allows it to effectively study flares in nearby active stars with observations not possible with any facility to date. For this reason, we will analyse this science case in detail, as it is Mauve’s “flagship science”. At the same time, other science areas can benefit from the same type of observations (spectrophotometric variability across the mid-near UV and visible bands). These include the study of variability in young, active stars such as Herbig Ae/Be stars, the study of variability in (more evolved) Be stars, monitoring of active binaries and characterisation of blue straggler stars.
@@ 5.1 Flaring solar-type stars
Our own Sun is a very quiet star compared to similar-type stars in the Galaxy. The advent of white light monitoring space telescopes dedicated to exoplanetary transit searches (e.g., Corot, Kepler, TESS) has shown that even in visible light many stars emit significant flares, much larger than anything observed in the Sun. However, during a flaring event, each wavelength probes a distinct region of the solar atmosphere. Consequently, while the white light curves provided by Kepler and TESS enable valuable statistical analyses of flare frequency across large stellar samples, they offer only limited potential for exploring the underlying physics of these events.
@@
A broader wavelength coverage of flaring events achievable with Mauve can enable the different spectral components to be disentangled. While with e.g. Kepler and TESS data, one sees the tail of the emission in white light, observations with Mauve would allow to test models of the emission from the heated stellar photosphere and lower chromosphere; different models of this exist, which scientists are interested to compare against actual observations. In particular, flares observed in white light are probably the result of light emitted from the heated, optically thick photosphere and lower chromosphere, that are likely to reach temperatures of 10,000 K or so, peaking in the near-UV band. By observing only the white light, one effectively sees the peak of the iceberg, while missing the bulk of it. With its wide wavelength coverage and the capacity to monitor specific sources for long periods, Mauve is ideal for studying individual large flares in detail, advancing the understanding of their physical mechanisms. Many exoplanets have been detected around active stars, and the UV emission from the host star dramatically influences the planetary evolution. Understanding the nature of this UV emission is crucial to gaining insights into broader topics, such as habitability.
@@
We know from Kepler and TESS data that some stars similar to the Sun undergo “super flares”, events in which their white light increases by several percent (events up to some 25% flux increase in the Kepler band have been observed, e.g. Shibayama et al. (2013) for the star KIC 12354328). If this flux rise is indeed produced by a heated patch of the photosphere, the increase in the blue and even more in the UV will be much larger, creating a “sweet spot” for Mauve observations. To demonstrate Mauve’s capabilities under different stellar brightness scenarios, we selected two example stars and used MauveSim to generate simulated observations. The example stars are a bright K2V star with $V_{\text{mag}}$=3.7 and a G2V star with a visual magnitude of 7.4.
5.1.1 K2V star
@@
Figure 10: Example input spectra of a K2V star at different flaring stages. The quiescent spectrum is displayed in a dashed blue line. The flaring spectra contain a blackbody component with a temperature of 10,000 K that becomes increasingly more pronounced as the flare approaches its peak.
@@
Figure 11: Simulated spectra of a K2V star at different flaring stages, as observed by Mauve. These spectra, generated with MauveSim, have been background-subtracted, i.e. the average noise floor has been removed from each data point and the data divided by the exposure time.
@@
The input spectra used to simulate the flare on the K2V star are shown in Fig. 10. These display the stellar spectrum at different flaring stages, from the quiescent phase indicated with a dashed blue line, to the flare peak in magenta, and at varying decaying stages (the rest of the spectra in between). The flaring spectra include an additional component to them, which is a simplistic model of a blackbody at 10,000 K that becomes increasingly stronger as we approach the peak flaring stage. By selecting an integration time of 50 s, we simulated how Mauve would see such spectra assuming the current performance. The corresponding MauveSim simulated spectra are shown in Fig. 11. The low resolution of the instrument does not allow to retrieve individual lines. However, the quiescence, flare peak and decay stages are clearly distinguishable from one another. We can clearly see a high contrast among spectra across the wavelength range accessible to Mauve, and specifically in the UV. Separating the various flare phases will be paramount for scientists who will want to test models of flare physics in solar-type stars.
@@ 5.1.2 G2V star
Figure 13: Simulated spectra of a G2V star at different flaring stages, as observed by Mauve. These spectra, generated with MauveSim, have been background-subtracted, i.e. the average noise floor has been removed from each data point and the data divided by the exposure time. Additionally, a Gaussian kernel with a FWHM of 10.5 nm was applied to the spectra to reduce noise and help distinguish the different flaring phases. This is an example near the limiting magnitude accessible to the Mauve’s flaring stars science case.
For bright targets (V magnitude brighter than 6) Mauve can do excellent science on large flares, even in the presence of the high level of dark current which is its limiting factor. When going to fainter magnitudes, the dark current becomes the limiting factor, and, at magnitudes of 7.5 or so, even for a large flare such one with a 25% increase in the Kepler band, the noise level will limit the science that can be achieved. It is worth recalling that while in a photon noise limited regime (where the source flux is low but the background is lower) integrating for a long time allows to achieve higher signal to noise observations (increasing with the square root of the observing time), in a background limited regime little if anything is gained through a long integration (as the background will also correspondingly increase).
Figure 13 provides an example that is considered to be near the limiting magnitude for Mauve’s flaring stars science case, i.e. the same large flare as for the K2V star, simulated for the fainter G2V stellar target in Fig. 12 ($V_{\text{mag}}$ = 7.4). While the spectrum at peak flare is clearly distinguishable from the quiescent phase, the noise level, in particular in the UV where the signal is low, is considerably high, so that it would be challenging to distinguish, e.g., different models for the flaring component.
@@ 6 Conclusions
At Blue Skies Space we developed MauveSim, an in-house instrument simulator to support the Mauve mission. This software generates simulated observations by requiring an input spectrum representative of the target source, which can either be an observed spectrum from other facilities or a synthetic model spectrum. The input spectrum is processed through the instrument response, incorporating its spectral resolution, optical element efficiencies, and detector response. Astrophysical and instrumental noise sources are then added to the net source signal. Users can specify the integration time, and the simulator produces a simulated raw spectrum as output. This tool is crucial in helping members of the Mauve science team refine their science cases and evaluate the feasibility of proposed observations. The software is accessible to all scientists who currently are or will be involved in the mission. Scientists not yet part of the Mauve survey are encouraged to explore the software by contacting BSSL by email at info@bssl.space.
@@ Acknowledgements
The Mauve project has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101082738.
@@
Data utilised as input to perform the simulations are publicly available in the HST/STIS Stellar Library at http://astro.wsu.edu/hststarlib/.

### v3

@@ 4.3 Converting flux into counts
$\text{efficiency}=GR\cdot FB\cdot(1-CL)^{4}\cdot MT^{2}\cdot MS^{2}\cdot DCL\cdot WD\cdot(1-LL)\,.$ (12)
@@
We combine the result thus obtained with the telescope’s collecting area $A$, the wavelength range per pixel $\Delta x_{\lambda}$, and the exposure time $t_{\text{exp}}$, which represents the duration of the observation. The input flux (in units of W/cm2 nm) is combined with the factors described above to convert the data in units of [A·s]. Subsequently, a series of conversion factors are applied to transform the data in units of counts. The conversion assumes that the 16-bit dynamic range is matched to the voltage dynamic range in the detector readout. According to the detector data sheet, the optimal voltage range is approximately $2.2\,\text{V}$. The total counts are calculated as:
@@ 4.4 Adding astrophysical and instrumental noise sources
Figure 8: Simulation of HD 37202 using an exposure time of 5 s. The final simulated observation and its components are shown as well. Panel (a): output Mauve spectrum as simulated by the software. Panel (b): net simulated signal source. Panel (c): photon noise. Panel (d): bias level. Panel (e): read noise.
Figure 9: Diagrams of the noise budget illustrating the contributions from each payload noise source, presented for different target types and at two distinct exposure durations: 50 s and 600 s. We considered the Be star HD 37202 and the K3V star HD 32147 which are bright targets for the mission, and HD 146233, a fainter G2V star. The plots show the source photon noise in green, detector dark current in magenta, the read noise in yellow and the photon noise from the zodiacal background in blue.
@@
Photon noise, which follows Poisson statistics, is modelled as the square root of the net signal in units of $e^{-}$. To simulate the distribution of photons striking the detector, we include a random component to the photon noise which is drawn from a Gaussian distribution with $\mu$=0 and $\sigma$$\sqrt{\text{net signal}[e^{-}]}$. Finally, to align the photon noise with the units of the source signal, photon noise is converted from $e^{-}$ to counts.
@@
where $T$ is the temperature of the detector. Based on this equation, the dark current is 445 counts/s at 21°C as reported in Table 2. This noise component is added by drawing random samples from a Poisson distribution with expectation value equal to $I_{\text{d}}(T)$ multiplied by the exposure time.
The bias is an offset applied to the ADC (Analogue-to-Digital Converter) to prevent negative counts during readout. The bias was measured in the lab and found to have an average value of 600 counts per pixel across the detector (Fig. 8 (d)).
Finally, readout noise is incorporated into the simulation (Fig. 8 (e)). It was estimated as the standard deviation of the bias for each pixel, determined from a set of 556 bias frames.
The combination of the average dark current value and bias is what we define as the noise floor.
To illustrate how the various noise sources scale with target magnitude, spectral type and exposure time, Fig. 9 shows the noise budget for the Be star HD 37202 and the flaring solar-type stars discussed in Sec. 5.1.
The MauveSim user interface in Stardrive prompts the user to specify the exposure time for which the simulation should be generated. The interface also alerts the user if the selected settings (exposure time and source magnitude) result in detector saturation. The software outputs a one-dimensional raw data spectrum, as illustrated in Fig. 8 (a). Mauve is planned to be operated with a 5-second cadence, meaning that each observation will consist of a series of 5-second frames. If the user selects an exposure time that is not a multiple of 5 seconds, the software will automatically round the exposure time up to the nearest multiple of 5. In this frame-stacking configuration, the signal from each frame is combined linearly, while the noise is added in quadrature.
@@ 4.5 Signal-to-noise calculation
Figure 10: The S/N per pixel of the simulated data in Fig. 8 (a).
Figure 11: S/N per resolution element obtained on the simulation of HD 37202 using an exposure time of 5 s.
@@
where $S$ is the net signal in units of counts (Fig. 8 (b)); $\sigma_{S}$ is the photon noise ($\sqrt{S}$); $\sigma_{R}$ is the readout noise (equal to approximately 12 counts); $\sigma_{D}$ is the dark current noise (equal to the sqrt of the dark current, i.e. $\sqrt{\sim 2220\ \text{counts}}$ for the example in Fig. 8). Fig. 10 displays the S/N per pixel of HD 37202 for a 5 s integration.
@@ 5 Science case studies
Mauve’s performance allows to effectively study flares in nearby active stars with observations not possible with any facility to date. For this reason, we will analyse this science case in detail, as it is Mauve’s “flagship science”. At the same time, other science areas can benefit from the same type of observations (spectrophotometric variability across the mid-near UV and visible bands). These include the study of variability in young, active stars such as Herbig Ae/Be stars, the study of variability in (more evolved) Be stars, monitoring of active binaries and characterisation of blue straggler stars.
@@ 5.1 Flaring solar-type stars
Our own Sun is a very quiet star compared to similar-type stars in the Galaxy. The advent of white light monitoring space telescopes dedicated to exoplanetary transit searches (e.g., Corot, Kepler, TESS) has shown that even in visible light many stars emit significant flares, much larger than anything observed in the Sun (Maehara et al., 2012). However, during a flaring event, each wavelength probes a distinct region of the solar atmosphere (Kowalski et al., 2019). Consequently, while the white light curves provided by Kepler and TESS enable valuable statistical analyses of flare frequency across large stellar samples, they offer only limited potential for exploring the underlying physics of these events (Hawley et al., 2014; Davenport, 2016; Günther et al., 2020).
@@
A broader wavelength coverage of flaring events achievable with Mauve can enable the different spectral components to be disentangled. While with e.g. Kepler and TESS data, one sees the tail of the emission in white light, observations with Mauve would allow to test models of the emission from the heated stellar photosphere and lower chromosphere; different models of this exist, which scientists are interested to compare against actual observations (Kowalski, 2024). In particular, flares observed in white light are probably the result of light emitted from the heated, optically thick photosphere and lower chromosphere, that are likely to reach temperatures of 10,000 K or so, peaking in the near-UV band (Kowalski et al., 2013). By observing only the white light, one effectively sees the peak of the iceberg, while missing the bulk of it. With its wide wavelength coverage and the capacity to monitor specific sources for long periods, Mauve is ideal for studying individual large flares in detail, advancing the understanding of their physical mechanisms. Many exoplanets have been detected around active stars, and the UV emission from the host star dramatically influences the planetary evolution (Segura et al., 2010; Segura, 2018; Howard et al., 2020). Understanding the nature of this UV emission is crucial to gaining insights into broader topics, such as habitability.
@@
We know from Kepler and TESS data that some stars similar to the Sun undergo “super flares”, events in which their white light increases by several percent (events up to some 25% flux increase in the Kepler band have been observed, e.g. Shibayama et al. (2013) for the star KIC 12354328). If this flux rise is indeed produced by a heated patch of the photosphere, the increase in the blue and even more in the UV will be much larger, creating a “sweet spot” for Mauve observations. To demonstrate Mauve’s capabilities under different stellar brightness scenarios, we selected two example stars and used MauveSim to generate simulated observations. The example stars are a bright K3V star with $V_{\text{mag}}$=3.7 and a G2V star with a visual magnitude of 7.4.
5.1.1 K3V star
@@
Figure 12: Example input spectra of a K3V star at different flaring stages. The quiescent spectrum is displayed in a dashed blue line. The flaring spectra contain a blackbody component with a temperature of 10,000 K that becomes increasingly more pronounced as the flare approaches its peak.
@@
Figure 13: Simulated spectra in units of counts per second of a K3V star at different flaring stages, as observed by Mauve. These spectra, generated with MauveSim, have been background-subtracted, i.e. the average noise floor has been removed from each data point and the resulting values divided by the exposure time.
@@
The input spectra used to simulate the flare on the K3V star are shown in Fig. 12. These display the stellar spectrum at different flaring stages, from the quiescent phase indicated with a dashed blue line, to the flare peak in magenta, and at varying decaying stages (the rest of the spectra in between). The flaring spectra include an additional component to them, which is a simplistic model of a blackbody at 10,000 K that becomes increasingly stronger as we approach the peak flaring stage. By selecting an exposure time of 50 s, we simulated how Mauve would see such spectra assuming the current performance. The corresponding MauveSim simulated spectra are shown in Fig. 13. The low resolution of the instrument does not allow to retrieve individual lines. A comparison of the flaring spectra with the quiescent phase using chi-squared analysis reveals that the early flaring stage and the decaying phase (0.76% and 2% flux increase) are statistically indistinguishable from the quiescent spectrum ($\bar{\chi}^{2}$ $\ll$ 1). Flare peak and mid-flare spectra (i.e. 24% and 8.1% flux increase) show strong, statistically significant deviations ($\bar{\chi}^{2}$ $\gg$ 1), particularly the flare peak, which returns the highest $\chi^{2}$ value. While the greatest spectral contrast is observed in the ultraviolet region, the capability to distinguish the various flaring phases will be crucial for scientists aiming to test models of flare physics in solar-type stars.
@@ 5.1.2 G2V star
Figure 15: Simulated spectra of a G2V star at different flaring stages, as observed by Mauve. These spectra, generated with MauveSim, have been background-subtracted, i.e. the average noise floor has been removed from each data point and the resulting values divided by the exposure time. Additionally, a Gaussian kernel with a FWHM of 10.5 nm was applied to the spectra to reduce noise and help to visually distinguish the different flaring phases. This is an example near the limiting magnitude accessible to the Mauve’s flaring stars science case.
For bright targets (V magnitude brighter than 6) Mauve can do excellent science on large flares, even in the presence of the high level of dark current. When going to fainter magnitudes, the dark current becomes the limiting factor, and, at magnitudes of 7.5 or so, even for a large flare such as one with a 25% increase in the Kepler band, the noise level will limit the science that can be achieved.
To demonstrate this limitation, Figure 15 provides an example that is considered to be near the limiting magnitude for Mauve’s flaring stars science case, i.e. the same large flare as for the K3V star, simulated for the fainter G2V stellar target in Fig. 14 ($V_{\text{mag}}$ = 7.4). A chi-squared comparison analysis between flaring and quiescent spectra reveals that the majority of the flaring spectra at the very beginning and end of the event are statistically indistinguishable from the quiescent phase. However, the flare peak stage creating a 24% flux increase from quiescence returns a high chi-squared value ($\chi^{2}$ = 6159.88, $\bar{\chi}^{2}$ = 3.75), indicating a significant deviation from the quiescent state. Similarly, the flaring stage with a 8.1% flux increase displays a moderate discrepancy from the quiescent phase ($\chi^{2}$ = 1078.36, $\bar{\chi}^{2}$ = 0.66). For a stellar target of this magnitude, it would be possible to distinguish the flare peak from quiescence. However, identifying a flaring event in the time series may be challenging unless the event is particularly energetic.
@@ 6 Conclusions
At Blue Skies Space we developed MauveSim, an in-house instrument simulator to support the Mauve mission. This software generates simulated observations by requiring an input spectrum representative of the target source, which can either be an observed spectrum from other facilities or a synthetic model spectrum. The input spectrum is processed through the instrument response, incorporating its spectral resolution, optical element efficiencies, and detector response. Astrophysical and instrumental noise sources are then added to the net source signal. Users can specify the exposure time, and the software produces a simulated raw spectrum as output. This tool is crucial in helping members of the Mauve science team refine their science cases and evaluate the feasibility of proposed observations. The software is accessible to all scientists who currently are or will be involved in the mission. Scientists not yet part of the Mauve survey are encouraged to explore the software by contacting BSSL by email at info@bssl.space.
@@ Acknowledgements
The Mauve project has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101082738. We thank ISISPACE, MediaLario and C3S for their partnership with BSSL in constructing and delivering Mauve. We also thank the anonymous referees for providing helpful comments that enhanced the quality of the manuscript.
@@
Data utilised as input to perform the simulations are publicly available in the HST/STIS Stellar Spectral Library at http://astro.wsu.edu/hststarlib/.